OpenStudy (anonymous):

Solve each system of equations by substitution. -2x+3y=11 x=4y-3

OpenStudy (noseboy908):

first put the 2nd eqn. into standard form.

OpenStudy (anonymous):

i did that but i still couldn't get the right answer

OpenStudy (noseboy908):

that's because there is more to it. what did you get when you put the 2nd eqn. into standard form?

OpenStudy (anonymous):

-2x + 3y = 11 3y = 11+2x Now, we know x = 4y-3 Substituting x=4y-3 in 3y=11+2x 3y = 11 + 2 (4y-3) 3y = 11 + 8y - 6 3y = 8y + 5 3y - 8y = 5 -5y = 5 y = -1 Substituting y = -1 in x = 4y-3 x = 4(-1) - 3 x = -4-3 = -7